Name these organic compounds: structure name О CH — CH, — 0= CH CH,– CH; O= CH CH; - O= CH – OH - CO
Answer #1

O=CH-CH2-CH3 name is Propanal.

O=CH-CH3 name is Ethanal.

O=CH-OH name is Methanoic acid.